OpenStudy (anonymous):

{(-5, -15), (-3, -9), (0, 0) , (2, 6) , (4, 12)} Identify the rule for the function. The range is equal to four more than the domain. The range is equal to eight less than the domain. The range is equal to three more than the domain. The range is equal to three times the domain.

OpenStudy (anonymous):

Look who joined the club... population one

OpenStudy (charlotte123):

range = y ^_^ so, the range is -15, -9, 0, 6, 12 ^_^ the domain is x, so the domain is -5, -3, 0, 2, 4 ^_^ think u can answer the question?

OpenStudy (anonymous):

Probably not >_<

OpenStudy (charlotte123):

hmm.......lets see

OpenStudy (anonymous):

this is my last question on my finals for the half the year test lol

OpenStudy (charlotte123):

The range is equal to three times the domain. ^_^ the other choices are >.>

OpenStudy (charlotte123):

-5 * 3 = -15, -3 * 3 = -9, and so on ^_^

OpenStudy (charlotte123):

makes sense?

OpenStudy (anonymous):

brb

OpenStudy (charlotte123):

okay ^_^ and welcome to OS :D

OpenStudy (anonymous):

just got back